### v1

# thanks: https://orcid.org/0009-0006-0705-752X

thanks: https://orcid.org/0009-0006-0705-752X
Towards fast DBSCAN via Spectrum-Preserving Data Compression
Yongyu Wang
preprint: APS/123-QED
This paper introduces a novel method to significantly accelerate DBSCAN by employing spectral data compression. The proposed approach reduces the size of the data set by a factor of five while preserving the essential clustering characteristics through an innovative spectral compression technique. This enables DBSCAN to run substantially faster without any loss of accuracy. Experiments on real-world data sets, such as USPS, demonstrate the method’s capability to achieve this dramatic reduction in data size while maintaining clustering performance.
I Introduction
Density-Based Spatial Clustering of Applications with Noise (DBSCAN) [1] has become a prominent method in the field of clustering, especially for identifying clusters of varying shapes and densities. Unlike traditional methods such as $k$-means, which assume spherical clusters and require the number of clusters to be specified beforehand, DBSCAN identifies clusters based on the density of data points, allowing it to effectively handle clusters of arbitrary shapes. Additionally, DBSCAN has the ability to distinguish noise from core cluster points, making it particularly suitable for real-world data with outliers. Compared to support vector clustering (SVC), DBSCAN does not rely on high-dimensional mappings or support vectors, but rather operates directly in the input space, making it more suitable density-based scenarios.
DBSCAN has gained popularity due to its ability to identify clusters of arbitrary shapes and handle noise effectively, making it well-suited for real-world applications. However, the original DBSCAN algorithm faces significant efficiency and scalability challenges when applied to large-scale datasets. Specifically, the neighborhood search step of DBSCAN has a $O(N^{2})$ time complexity [2], as it requires pairwise distance calculations to identify core points within a specified radius ($\varepsilon$). This quadratic complexity can become prohibitive in big data environments, where the time required for clustering grows rapidly with the number of data points. Additionally, DBSCAN’s performance degrades with high-dimensional data due to the curse of dimensionality, where distance metrics become less meaningful, further impacting both speed and clustering quality. These limitations restrict DBSCAN’s applicability in emerging big data analytics tasks, necessitating an optimized approach to improve its scalability and efficiency.
Inpsired by [3], in this paper, we propose to fundamentally address the efficiency and scalability issues of DBSCAN through a novel spectrum-preserving data compression method. Our approach first computes a significantly smaller, compressed data set while preserving the essential clustering properties of the original data, utilizing a nearly-linear time spectral data compression technique. This compressed data set allows us to dramatically reduce the search space for density-based clustering, making DBSCAN feasible for large-scale applications. The compressed data then serves as input for an accelerated DBSCAN algorithm, where the original clustering structure is retained, not only without loss in solution quality. An additional benefit of the proposed method is that users can specify the compression ratio based on their requirements.
II Preliminary
II.1 DBSCAN
Given a dataset $D$ with $N$ samples $\mathbf{x}_{1},\dots,\mathbf{x}_{N}$, the DBSCAN (Density-Based Spatial Clustering of Applications with Noise) algorithm identifies clusters through two main steps: neighborhood search and cluster expansion.
In the neighborhood search step, DBSCAN defines clusters based on the density of data points. For a given point $\mathbf{x}_{i}$, the algorithm examines all points within a specified radius $\varepsilon$ (referred to as the $\varepsilon$-neighborhood). A point $\mathbf{x}_{i}$ is classified as a core point if the number of points within its $\varepsilon$-neighborhood, including $\mathbf{x}_{i}$ itself, is at least a minimum threshold MinPts. Points that do not meet this criterion are classified either as border points (if they lie within the $\varepsilon$-neighborhood of a core point) or as noise if they do not belong to any cluster.
Once core points are identified, DBSCAN proceeds to the cluster expansion step. Here, each core point initiates a cluster by linking to other core points within its $\varepsilon$-neighborhood, iteratively expanding the cluster by including reachable core points and their neighbors. This expansion continues until all reachable points within the density threshold are included in the cluster. The process is repeated for all core points, eventually identifying all clusters in the dataset.
Formally, DBSCAN can be summarized by the following steps:
  1. 1.
    Identify Core Points: For each point $\mathbf{x}_{i}$, compute the $\varepsilon$-neighborhood and count the number of points within it. If the count meets or exceeds MinPts, classify $\mathbf{x}_{i}$ as a core point.
  2. 2.
    Cluster Expansion: For each core point, form a new cluster and iteratively add all directly reachable core points within $\varepsilon$-neighborhoods. Expand the cluster until no more points meet the density-reachability criteria.
  3. 3.
    Classify Border and Noise Points: Points within the $\varepsilon$-neighborhood of core points but with fewer than MinPts neighbors are classified as border points. Points that do not fall within any cluster are classified as noise.
DBSCAN’s main advantage lies in its ability to detect clusters of arbitrary shape and to identify noise. However, it also has limitations, particularly in terms of computational complexity. The algorithm’s neighborhood search step has a time complexity of $\mathbf{O(N^{2})}$, as it involves pairwise distance calculations for each point. This quadratic complexity can become prohibitive for large-scale datasets, especially in high-dimensional spaces where distance calculations are computationally intensive.
II.2 Graph Laplacian Matrices
Consider a graph $G=(V,E,w)$, where $V$ represents the set of vertices, $E$ is the set of edges, and $w$ is a function assigning positive weights to each edge. The Laplacian matrix of graph $G$, which is a symmetric diagonally dominant (SDD) matrix, is defined as follows:
$L_{G}(p,q)=\begin{cases}-w(p,q)&\text{if }(p,q)\in E,\\ \sum\limits_{(p,t)\in E}w(p,t)&\text{if }p=q,\\ 0&\text{otherwise}.\end{cases}$ (1)
According to spectral graph theory [4], the spectrum of the graph Laplacian contains valuable information about the clustering structure of the graph, which is essential for identifying clusters. For instance, in artificially generated datasets such as two-moons and two-circles, the clustering results of $k$-means applied directly in the original feature space differ significantly from those in the spectrally-embedded feature space. In the original space, the traditional $k$-means algorithm is unable to correctly identify the clusters due to its reliance on Euclidean distance alone. However, by incorporating spectral information into the feature space, the clustering becomes accurate and aligns with the true underlying structure.
III Methods
III.1 Algorithmic Framework
Given a dataset of samples, we initially build a $k$-nearest neighbors (k-NN) graph $G$, which captures the fundamental structure or manifold of the dataset. The corresponding graph Laplacian matrix $L_{G}$ can then be derived from $G$. The spectral similarity between any two points in the dataset is calculated through the following steps:
  • $\bullet$
    Compute the bottom $K$ eigenvectors of the graph Laplacian that are nontrivial;
  • $\bullet$
    Form a matrix $U$ by arranging these $K$ eigenvectors as column vectors;
  • $\bullet$
    Each row of $U$ serves as a low-dimensional embedded feature vector for a point;
  • $\bullet$
    Define the spectral similarity between two points, $u$ and $v$, as follows [5]:
$s_{uv}:=\frac{{|(\mathbf{X}_{u},\mathbf{X}_{v})|}^{2}}{(\mathbf{X}_{u},\mathbf{X}_{u})(\mathbf{X}_{v},\mathbf{X}_{v})},\quad(\mathbf{X}_{u},\mathbf{X}_{v}):=\sum_{k=1}^{K}{\left(\mathbf{x}_{u}^{(k)}\cdot\mathbf{x}_{v}^{(k)}\right)}.$ (2)
Here, $\mathbf{x}_{u}$ and $\mathbf{x}_{v}$ are the feature vectors of points $u$ and $v$ within the embedded space. Interpreted statistically, $s_{uv}$ represents the fraction of variance explained by the linear regression of $\mathbf{x}_{u}$ on $\mathbf{x}_{v}$ from a spectral viewpoint. A larger value of $s_{uv}$ implies a stronger spectral connection between the points. To reduce the computational burden associated with eigen-decomposition, Gauss-Seidel relaxation can be employed to produce smoothed vectors for embedding features, achieving linear-time complexity [5].
By using spectral similarity, we divide the samples into subsets where samples in the same subset exhibit high spectral correlation. The mean of feature vectors within each subset is calculated to create a spectrally representative pseudo-sample, and these pseudo-samples collectively form a compressed version of the dataset. If further compression is needed, these representative pseudo-samples can be compressed again. Fig. 1 and Fig. 2 illustrate the manifolds of the original USPS dataset and the compressed dataset, respectively, by projecting their $k$-NN graphs in 2D. This comparison reveals that the compressed data effectively approximates the original manifold structure.
Figure 1: The $k$-NN graph corresponding to the original data set (USPS).
Figure 2: The $k$-NN graph corresponding to the compressed data set (USPS).
Then, the DBSCAN process can be applied on the compressed data points to group them into distinct clusters. Finally, each original data point is assigned the same cluster membership as its corresponding compressed point, completing the clustering task. The entire algorithm flow is illustrated as follows:
Input: A data set $D$ with $N$ samples $x_{1},\dots,x_{N}\in\mathbb{R}^{d}$.
Output: Clusters $C_{1},\dots,C_{k}$.
1:  Construct a $k$-nearest neighbor ($k$-NN) graph $G$ from the input data.
2:  Compute the Laplacian matrix corresponding to graph $G$.
3:  Perform spectrum-preserving data compression on the original data set.
4:  Perform DBSCAN on the compressed data set.
5:  Map the cluster memberships of compressed points back to obtain the clustering result of the original data set.
III.2 Algorithm Complexity
The spectrum-preserving data compression requires $O(|E_{G}|\log(|V|))$ time, where $|E_{G}|$ denotes the number of edges in the original graph, and $|V|$ represents the number of data points. Applying DBSCAN on the compressed dataset has a complexity of $O(P^{2})$, where $P$ is the number of compressed points. Retrieving the cluster memberships for the original data points has a time complexity of $O(N)$, where $N$ is the total number of samples in the original dataset.
IV Experiment
We use USPS data set to evaluate the method. NMI [6] is used as the metrics.
Orig DBSCAN Ours(5X reduction)
0.5513 0.5517
Table 1: Clustering Quality Results (NMI)
V Conclusion
To address the computational challenges of DBSCAN, this study presents a novel spectral data compression approach that creates a significantly smaller data set while preserving the essential spectral properties needed for clustering. Experiments on large-scale, real-world data sets demonstrate that the proposed method can significantly accelerates DBSCAN algorithm.